OpenStudy (anonymous):

hurrry hurrry please help got a test on a time limit 4/x+4=6/x (their like set up as in faction form )

OpenStudy (turingtest):

multiply the entire expression y x what do you get?

OpenStudy (turingtest):

sorry *by x

OpenStudy (turingtest):

\[x\cdot(\frac4x+4)=x\cdot(\frac6x)\]

OpenStudy (turingtest):

what will be left?

OpenStudy (anonymous):

1/2?